OpenStudy (australopithecus):

Anyone able to explain DeMovire's Theorem? If someone could walk me through this problem it would be greatly appreciated: Find all fourth roots of \[8(1 + \sqrt{3}i)\]; that is, find all complex numbers z such that \[z^{4} = 8(1 + \sqrt{3}i)\] Polar coordinate \[r = \sqrt{1^2 + (\sqrt3)^2} = 2\] \[\theta = \arcsin(\frac{1}{2}) = \frac{\pi}{3}\] thus we have, \[z=8(2e^{\frac{\pi}{3}i})\] The final answer is, z_0 = 2e^pi(i)/12 z_1 =2e^(7pi)i/12 z_2 = 2e^(13pi)i/12 z_3 = 2e^(19pi)i/12

OpenStudy (loser66):

take a look at http://en.wikipedia.org/wiki/De_Moivre%27s_formula

OpenStudy (australopithecus):

DeMovire's Theorem is simply \[z^n = (r(\cos(\theta) + \sin(\theta)i)^n\] z^n = (r(cos(theta) + sin(theta)i)^n\[z^n = r^n(\cos(\theta n) + \sin(\theta n)i)\] Then you can easily convert back to standard form. No one will probably look at this but smeh